OpenStudy (anonymous):

find the general form of the equation of the circle. Center at the origin and containing the point (-2,3)

OpenStudy (anonymous):

Generally, the formula for any circle centred on the point \((a,b)\) is: \[(x-a)^2+(y-b)^2=r^2\] So if it is centred on the origin, you have \[x^2+y^2=r^2\] And if you know a point, you can find out the radius by plugging in the x and y co-ordinates to find your radius and therefore complete the equation: \[(-2)^2+(3)^2=r^2\] \[4+9=r^2\] \[13=r^2\] \[r=\sqrt{13}\] So therefore, the formula is \[x^2+y^2=13\]

OpenStudy (anonymous):

thanks!!!!!

OpenStudy (anonymous):

Anytime man. You can also solve for x by doing: \[x^2+y^2=13\] \[y^2=13-x^2\] \[y=\pm\sqrt{13-x^2}\]